# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

### v2

@@ F.1 Evaluation of PET on Movielens
Table 3: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B. Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold.
@@
Table 4: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold.
@@
Table 5: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold.
F.2 Evaluation of PET on Yelp
@@
Table 6: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold. Level: L1
@@
Table 7: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold. Level: L1
@@
Table 8: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold. Level: L1

### v3

@@ F.1 Evaluation of SPECTRA on Movielens
Table 7: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B. Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold.
@@
Table 8: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold.
@@
Table 9: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold.
(a) Group-level preference evolution as a probability 3D map
(b) Group-level preference evolution as a bar chart
Figure 5: Evolution of group-level movie genre preferences as probability distribution across four time periods. Dataset: Movielens, 125 users over 4 periods. Method: SPECTRA Likelihood-based Probing with CPT+SFT on Qwen3-8B. Note: due to limited space, we only show top-10 clusters (genres) here.
F.2 Evaluation of SPECTRA on Yelp
Qwen3-8B DeepSeek-distill-Llama-8B
Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
Logit Probing .980 .879 .915 .925 .856 .880
Generative Classification .107 .118 .163 .001 .062 .119
L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
Logit Probing .513 .562 .666 .261 .268 .363
Generative Classification .466 .349 .454 .684 .553 .574
Table 10: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task. Results are shown for a context window of 8 using the CPT+SFT pipeline. Best NDCG scores for each level are in bold. L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”. Alternative taxonomies can be easily integrated depending on platform objectives. Sample Size: 5000.
@@ F.2 Evaluation of SPECTRA on Yelp
Table 11: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold. Level: L1
@@
Table 12: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold. Level: L1
@@
Table 13: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold. Level: L1